The solution set for the system 6x + 9y = 15 is best described as follows:
4x + 6y = 10
a.undefined
b.x=1. y=1
c. {(x,y)|y= (-2/3)x + (5/3)}
d. inconsistent
e. set of all points

Answer by stanbon(75887)

The solution set for the system
6x + 9y = 15
4x + 6y = 10
is best described as follows:
-----
slope of the 1st is -2/3 ; intercept = 5/3
slope of the 2nd is -2/3 ; intercept = 5/3
-----
Ans: The two equations are the same line.
The answer is "c".
=======================
Cheers,
Stan H.
====================
a.undefined
b.x=1. y=1
c. {(x,y)|y= (-2/3)x + (5/3)}
d. inconsistent
e. set of all points